# CoA: Chain-of-Action for Generative Semantic Labels

### v1

@@ 3.3 Chain-of-Action
Based on the analysis in the previous section, VQA-based LMMs struggle to produce ideal labels due to significant semantic overlap, while Caption-based LMMs face challenges with high parameter tuning costs, poor adaptability, and heavy reliance on filtering strategies. However, we observed that each interaction with LMMs provides valuable contextual information that can enhance the identification of target entities. Inspired by the success of Chain-of-Thought techniques in large language models and the application of In-context learning in multimodal tasks [15, 32, 20, 37], we propose a simple yet effective Chain-of-Action approach for the GSLs task.
Method Split-0 Split-1 Split-2 Split-3 Avg
Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$
VQA BLIP-2 [24] 46.00 68.83 32.39 69.10 41.35 66.40 25.00 70.72 36.19 68.76
InstructBLIP [9] 36.45 72.14 33.97 72.93 42.57 69.12 21.14 68.34 33.53 70.63
LLaVA [29] 55.22 66.50 53.16 62.05 54.99 66.66 55.87 66.90 54.81 65.53
MiniGPT-4 [58] 45.87 45.64 52.17 45.32 54.73 44.76 49.16 48.48 50.48 46.05
Caption BLIP-2 [24] 64.45 69.87 55.27 70.03 66.58 72.57 36.74 69.12 55.76 70.40
InstructBLIP [9] 69.37 69.12 69.69 70.54 73.44 72.08 71.04 74.05 70.89 71.48
LLaVA [29] 58.88 60.16 66.06 65.28 63.62 67.93 67.21 65.57 63.94 64.74
MiniGPT-4 [58] 46.76 40.34 52.98 37.32 52.09 39.62 45.30 43.63 49.28 40.23
@@
Method Split-0 Split-1 Split-2 Split-3 Avg
Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$
VQA BLIP-2 [24] 34.19 68.23 25.40 62.67 42.03 72.39 37.61 70.25 34.81 68.39
InstructBLIP [9] 34.14 70.14 26.35 59.98 43.87 72.47 31.36 70.89 33.93 68.37
LLaVA [29] 54.16 65.85 47.28 65.09 70.14 69.04 55.87 72.32 56.86 68.08
MiniGPT-4 [58] 38.52 50.48 40.88 55.98 44.44 55.25 57.16 60.04 45.25 55.44
Caption BLIP-2 [24] 60.17 62.35 41.00 65.65 62.79 64.76 58.70 72.66 55.67 66.36
InstructBLIP [9] 64.73 75.20 61.08 72.78 61.91 77.44 63.00 71.71 62.68 74.28
LLaVA [29] 67.67 63.29 67.63 66.84 75.42 75.96 69.61 71.70 70.08 69.45
MiniGPT-4 [58] 44.28 45.26 39.60 49.69 47.66 51.60 58.99 51.52 47.63 49.52
@@
[…]
Method Split-0 Split-1 Split-2 Split-3 Avg
Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$
VQA BLIP-2 [24] 70.76 60.54 70.52 59.94 71.88 59.99 71.48 60.74 71.16 60.30
InstructBLIP [9] 64.40 62.70 64.96 62.07 65.20 62.13 66.16 62.52 65.18 62.36
LLaVA [29] 71.68 60.33 71.00 61.27 73.12 59.80 71.12 60.62 71.73 60.51
MiniGPT-4 [58] 76.76 49.97 77.32 50.30 70.36 49.15 69.84 50.30 73.57 49.93
Caption BLIP-2 [24] 72.68 63.01 70.92 61.98 73.08 62.31 73.36 62.56 72.51 62.47
InstructBLIP [9] 72.24 65.73 72.88 66.15 72.76 64.98 72.12 65.57 72.50 65.61
LLaVA [29] 74.76 61.71 73.32 63.77 75.08 61.21 72.44 63.80 73.90 62.62
MiniGPT-4 [58] 77.84 43.62 78.00 43.98 73.80 44.40 71.36 44.06 75.25 44.02
@@
[…]
We define each interaction with LMMs as an individual action. By employing different combinations of prompts (as illustrated in Fig. 4), we interact with the LMMs to obtain complementary outputs. We establish a five-step process to guide the LMMs in producing: (i) Caption Action: Generate a one-sentence caption for the image and use a filtering strategy (detailed in the Appendix B) to extract an initial list of entities. (ii) Utilize Self-Correct Action to perform self-correction on each entity in the initial object list (isolating single-entity corrections improves instruction specificity, as LMMs handle Yes-or-No questions more effectively than open-ended What-and-How questions). (iii) Apply Appearance Action to extract each entity’s appearance and attributes. (iv) Using Relationship Action, infer relationships between entities by combining the corrected entity list with extracted appearance details. (v) Employ Final Action to integrate the entity list with outputs from Appearance Action and Relationship Action, generating the final entity list. To provide a comprehensive understanding of the proposed method, Fig. 3 illustrates the architecture and execution pipeline.
The proposed CoA method is capable of generating all relevant entity labels for an image without relying on predefined label sets or external database retrieval. The Caption Action step mitigates the issue of generating semantically redundant labels, while the Self-Correct Action ensures self-verification of the filtered labels, further unlocking the potential of LMMs. Additionally, the Appearance Action enhances the capture of local entity details, and the Relationship Action enriches the representation of inter-entity relationships, providing contextual depth for the Final Action. This progressive structure ensures that the LMMs’ output maintains high accuracy while improving the number of predicted entities. These findings are validated through extensive experiments, as detailed in Section 4.
Dataset ##\## Split-0 ##\## Split-1 ##\## Split-2 ##\## Split-3 ##\## Total
VOC 1561 1775 1891 596 5823
COCO 29628 3583 4461 2465 40137
NUS 2500 2500 2500 2500 10000
Table 3: The details of datatsets.
4 Experiment
Type Split-0 Split-1 Split-2 Split-3 Avg
Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$
(i) LLaVA w/Action 5 67.65 74.64 67.38 76.89 75.15 75.54 65.77 74.27 68.99 75.34
@@ 4 Experiment
[…]
Type Split-0 Split-1 Split-2 Split-3 Avg
Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$
LLaVA-1.5-7B 55.22 66.50 53.16 62.05 54.99 66.66 55.87 66.9 54.81 65.53
LLaVA-1.5-7B w/ CoA 81.23 (+26.01) 75.18 (+8.68) 79.83 (+26.67) 78.54 (+16.49) 84.51 (+29.52) 79.69 (+13.03) 84.23 (+28.36) 76.54 (+9.64) 82.45 (+27.64) 77.49 (+11.96)
LLaVA-1.5-13B 68.74 69.67 68.34 67.77 77.95 66.39 68.62 68.5 70.91 68.08
LLaVA-1.5-13B w/ CoA 81.42 (+12.68) 76.94 (+7.27) 81.63 (+13.29) 79.72 (+11.95) 86.94 (+9.99) 84.35 (+17.96) 90.76 (+22.14) 79.70 (+11.20) 85.19 (+14.28) 80.18 (+12.10)
[…]
In this section, we present an extensive overview of the datasets, models, baselines, implementation details and evaluation metrics utilized in all experiments (see Section 4.1). Furthermore, we provide extensive quantitative and qualitative experiments in Sections 4.2 $\sim$ 4.4 to validate the efficacy of the proposed method and compare its effectiveness to the existing state-of-the-art methods.
4.1 Setup
Datasets. The core of the GSLs task lies in identifying multiple entities within an image. To accurately assess the effectiveness of our approach, we employ widely used multi-label datasets for evaluation, including PASCAL VOC 2012 (VOC) [12], MS-COCO 2014 (COCO) [26], and NUS-WIDE (NUS) [7]. As our method operates in a zero-shot inference setting without training, we use only the test sets from these datasets. Specifically, following [4, 52], we divided each dataset into four non-overlapping subsets. For the NUS dataset, we validate our method’s effectiveness on a test set created by randomly sampling 10,000 images from the original NUS test set. The more details of these datasets can be obtained from Table 3 and Appendix E.
Models. The pre-trained LLaVA-1.5-7B [29] model is used as the base model, and we apply our CoA approach to this model to address GSLs tasks. Furthermore, we compare our approach with three popular LMMs: BLIP-2 [24], InstructBLIP [9], and MiniGPT-4 [58]. Additional details are provided in Appendix C.
Baselines. We categorize our baselines into two main groups for comparative analysis. The first group involves using VQA models to directly predict class names associated with an image. This includes BLIP-2 [24], InstructBLIP [9], LLaVA [29], and MiniGPT-4 [58], all of which have demonstrated outstanding performance in VQA tasks. The second group consists of captioning approaches, as captions effectively encapsulate the semantic content of images. For these, we also evaluate the same four LMMs, given their proven efficacy in caption generation.
Implementation Details. All experiments were conducted on two NVIDIA 4090 GPUs. Following CaSED [8] and the BLIP-2 [24] demo, for captioning, we used the prompt “Question: what’s in the image? Answer:”. For VQA, we used the prompt “Question: What are the names of objects in this image? Answer:”.
Evaluation metrics. Given that the GSLs task requires identifying multiple entities, we aim for model predictions to be as comprehensive and accurate as possible. Therefore, we designed two evaluation metrics that separately assess comprehensiveness ($\text{M}_{\text{clip}}$) and accuracy ($\text{M}_{\text{ram}}$). Due to page limitations, we provide details on the definitions and calculation formulas for these two metrics in Appendix D.
4.2 Quantitative results
[…]
Type Split-0 Split-1 Split-2 Split-3 Avg
Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$ Mclip $\uparrow$ Mram $\uparrow$
CoA 81.23 75.18 79.83 78.54 84.51 79.69 84.23 76.54 82.45 77.49
Single Interaction 65.86 (-15.37) 49.18 (-26.00) 62.42 (-17.41) 39.89 (-38.65) 71.44 (-13.07) 42.79 (-36.90) 67.28 (-16.95) 36.97 (-39.57) 66.75 (-15.70) 42.21 (-35.28)
[…]

### v2

@@ 3.3. Chain-of-Action
Based on the analysis in the previous section, VQA-based methods struggle to produce ideal labels due to significant semantic overlap, while Caption-based methods face challenges with high parameter tuning costs, poor adaptability, and heavy reliance on filtering strategies. However, we observed that each interaction with VLMs provides valuable contextual information that can enhance the identification of target entities. Inspired by the success of Chain-of-Thought techniques in large language models and the application of In-context learning in multimodal tasks (Gao and Das, 2024; Nie et al., 2024; Koike et al., [n. d.]; Sheng et al., 2024), we propose a simple yet effective Chain-of-Action approach for the GSLs task.
We define each interaction with VLMs as an individual action. By employing different combinations of prompts (as illustrated in Fig. 4), we interact with the VLMs to obtain complementary outputs. We establish a five-step process to guide VLMs in generating comprehensive and accurate semantic labels. Each step is designed to progressively refine the model’s understanding of the image, ensuring that the final output captures both the granular details and the broader contextual relationships within the scene.
(i) Caption Action. The process begins with generating a one-sentence caption for the image, which serves as a high-level summary of its content. From this caption, we apply a filtering strategy to extract an initial list of entities. The instruction used for the Caption Action is selected as the optimal prompt from a predefined set of caption instructions. This selection is based on optimization conducted over a small auxiliary dataset. This action provides a foundational understanding of the image, identifying key objects or concepts that will be further refined in subsequent steps.
(ii) Self-Correct Action. Next, we perform a self-correction on each entity in the initial list. This step isolates individual entities and leverages the VLM’s ability to handle Yes-or-No questions more effectively than open-ended queries. By asking targeted questions (e.g., ”Is there a [object] in the image?”), we ensure that each entity is accurately identified and remove any false positives or ambiguities. This refinement enhances the precision of the entity list, setting the stage for more detailed analysis.
(iii) Appearance Action. With a corrected entity list in hand, we proceed to extract appearance and attribute details for each entity. This step involves querying the VLM to describe the visual characteristics of each object (e.g., ”What color is the [object]?” or ”What material is the [object] made of?”). By enriching the entity list with descriptive attributes, we capture finer-grained semantic information that is crucial for understanding the image’s composition.
(iv) Relationship Action. Building on the enriched entity list, we infer relationships between entities by combining the corrected entity list with the extracted appearance details. This step involves asking the VLM to identify interactions or spatial relationships (e.g., ”Is the [object1] next to the [object2]?” or ”Is the [object1] holding the [object2]?”). By contextualizing entities within their relationships, we move beyond isolated object recognition to a more holistic understanding of the scene.
(v) Final Action. In the final step, we integrate the entity list, appearance details, and relationship inferences to generate a comprehensive set of semantic labels. This step synthesizes the outputs from the previous actions, ensuring that the final labels are not only accurate but also contextually rich. The result is a detailed and structured representation of the image’s semantic content, suitable for downstream tasks such as image retrieval, content generation, or scene understanding.
By breaking down the generative process into these five steps, our approach ensures that each action builds on the previous one, progressively refining the model’s understanding and enabling it to produce high-quality semantic labels. This structured methodology not only improves the accuracy of the generated labels but also enhances the interpretability and adaptability of the model in open-ended domains. To provide a comprehensive understanding of the proposed method, Fig. 2 illustrates the architecture and execution pipeline.
The proposed CoA method is capable of generating all relevant entity labels for an image without relying on predefined labels set or external database retrieval. The Caption Action step mitigates the issue of generating semantically redundant labels, while the Self-Correct Action ensures self-verification of the filtered labels, further unlocking the potential of VLMs. Additionally, the Appearance Action enhances the capture of local entity details, and the Relationship Action enriches the representation of inter-entity relationships, providing contextual depth for the Final Action. This progressive structure ensures that the VLMs’ output maintains high accuracy while improving the number of predicted entities. These findings are validated through extensive experiments in Section 4.
4. Experiment
In this section, we present an extensive overview of the datasets, models, baselines, implementation details and evaluation metrics utilized in all experiments (see Section 4.1). Furthermore, we provide extensive quantitative and qualitative experiments in Sections 4.2 $\sim$ 4.7 to validate the efficacy of the proposed method and compare its effectiveness to the existing state-of-the-art methods.
Table 1. Comparison of results between the proposed CoA and various VQA-based and Caption-based methods in $\text{M}_{\text{com}}$ ($\%$) and $\text{M}_{\text{acc}}$ ($\%$) metrics on the VOC and COCO datatsets. The best performance is highlighted in bold.
Method Vehicles Animals Indoor Outdoor Avg
Mcom $\uparrow$ Macc $\uparrow$ Mcom $\uparrow$ Macc $\uparrow$ Mcom $\uparrow$ Macc $\uparrow$ Mcom $\uparrow$ Macc $\uparrow$ Mcom $\uparrow$ Macc $\uparrow$
VQA BLIP-2 (Li et al., 2023) 46.00 68.83 32.39 69.10 41.35 66.40 25.00 70.72 36.19 68.76
InstructBLIP (Dai et al., 2023) 36.45 72.14 33.97 72.93 42.57 69.12 21.14 68.34 33.53 70.63
LLaVA (Liu et al., 2024c) 55.22 66.50 53.16 62.05 54.99 66.66 55.87 66.90 54.81 65.53
MiniGPT-4 (Zhu et al., 2024b) 45.87 45.64 52.17 45.32 54.73 44.76 49.16 48.48 50.48 46.05
Caption BLIP-2 (Li et al., 2023) 64.45 69.87 55.27 70.03 66.58 72.57 36.74 69.12 55.76 70.40
InstructBLIP (Dai et al., 2023) 69.37 69.12 69.69 70.54 73.44 72.08 71.04 74.05 70.89 71.48
LLaVA (Liu et al., 2024c) 67.65 74.64 67.38 76.89 75.15 75.54 65.77 74.27 68.99 75.34
MiniGPT-4 (Zhu et al., 2024b) 46.76 40.34 52.98 37.32 52.09 39.62 45.30 43.63 49.28 40.23
@@ 4. Experiment
Method Organism Vehicles Indoor Outdoor Avg
Mcom $\uparrow$ Macc $\uparrow$ Mcom $\uparrow$ Macc $\uparrow$ Mcom $\uparrow$ Macc $\uparrow$ Mcom $\uparrow$ Macc $\uparrow$ Mcom $\uparrow$ Macc $\uparrow$
VQA BLIP-2 (Li et al., 2023) 34.19 68.23 25.40 62.67 42.03 72.39 37.61 70.25 34.81 68.39
InstructBLIP (Dai et al., 2023) 34.14 70.14 26.35 59.98 43.87 72.47 31.36 70.89 33.93 68.37
LLaVA (Liu et al., 2024c) 54.16 65.85 47.28 65.09 70.14 69.04 55.87 72.32 56.86 68.08
MiniGPT-4 (Zhu et al., 2024b) 38.52 50.48 40.88 55.98 44.44 55.25 57.16 60.04 45.25 55.44
Caption BLIP-2 (Li et al., 2023) 60.17 62.35 41.00 65.65 62.79 64.76 58.70 72.66 55.67 66.36
InstructBLIP (Dai et al., 2023) 64.73 75.20 61.08 72.78 61.91 77.44 63.00 71.71 62.68 74.28
LLaVA (Liu et al., 2024c) 67.67 63.29 67.63 66.84 75.42 75.96 69.61 71.70 70.08 69.45
MiniGPT-4 (Zhu et al., 2024b) 44.28 45.26 39.60 49.69 47.66 51.60 58.99 51.52 47.63 49.52
@@
4.1. Experiment Setup
Datasets. The core of the GSLs task lies in identifying multiple entities within an image. To accurately assess the effectiveness of our approach, we employ widely used multi-label datasets for evaluation, including PASCAL VOC 2012 (VOC) (Everingham et al., 2012), MS-COCO 2014 (COCO) (Lin et al., 2014), and NUS-WIDE (NUS) (Chua et al., 2009). As our method operates in a zero-shot inference setting without training, we use only the test sets from these datasets. Specifically, following (Bar et al., 2022; Zhang et al., 2023b), we divided each dataset into four non-overlapping subsets. In addition, we collected a real-world open-vocabulary dataset (OVD) to evaluate our method, consisting of 200 test images spanning both indoor and outdoor scenarios.
Models. The pre-trained LLaVA-1.5-7B (Liu et al., 2024c) model is used as the base model, and we apply our CoA approach to this model to address GSLs tasks. Furthermore, we compare our approach with three popular VLMs: BLIP-2 (Li et al., 2023), InstructBLIP (Dai et al., 2023), and MiniGPT-4 (Zhu et al., 2024b).
Baselines. We categorize our baselines into two main groups for comparative analysis. The first group involves using VQA models to directly predict class names associated with an image. This includes BLIP-2 (Li et al., 2023), InstructBLIP (Dai et al., 2023), LLaVA (Liu et al., 2024c), and MiniGPT-4 (Zhu et al., 2024b), all of which have demonstrated outstanding performance in VQA tasks. The second group consists of captioning approaches, as captions effectively encapsulate the semantic content of images. For these, we also evaluate the same four VLMs, given their proven efficacy in caption generation. We utilize different sets of VQA and captioning instructions to obtain the optimal results for these baselines.
Evaluation Metrics. Evaluating the effectiveness of methods for GSLs is inherently challenging due to the unbounded and unconstrained nature of the semantic space. In this paper, we propose two main evaluation metrics namely Semantic Comprehensiveness ($\text{M}_{\text{com}}$), i.e. the overall similarity of the predicted labels w.r.t the given image, and Semantic Accuracy ($\text{M}_{\text{acc}}$), i.e. the quantity and quality of the predicted labels.
Implementation Details. All experiments were conducted on two NVIDIA 4090 GPUs. For Semantic Comprehensiveness, we used the prompt $P_{com}=$ ”This image contains […]” to capture the comprehensiveness of predicted entities. Next, given the predicted entity list $Y_{i}$ and the manually annotated entity list $\bar{Y}_{i}$, we compute their respective similarities to the image. Then, we can obtain the comprehensive score $CS_{i}$ or the predicted labels of each test image $I_{i}$ from the following formula:
$\small CS_{i}=\mathbf{1}\left[cos\langle T(P_{com}(Y_{i})),V(I_{i})\rangle>cos\langle T(P_{com}(\bar{Y}_{i})),V(I_{i})\rangle\right],$
where $\mathbf{1}[\cdot]$ denotes the indicator function, $T(\cdot)$ and $V(\cdot)$ denotes the text and image encoder of CLIP model. A prediction is scored only if the similarity of the model-generated prompt exceeds that of the prompt based on manual annotations. Here, we utilize the label set provided by the dataset itself as the manual-annotated labels. Then, we can obtain the $\text{M}_{\text{com}}=\sum_{i}^{N}CS_{i}/N$ to assess the coverage of predicted labels, where $N$ denotes the number of test images.
Table 2. Comparison of results between the proposed CoA and various VQA-based and Caption-based methods in $\text{M}_{\text{com}}$ ($\%$) and $\text{M}_{\text{acc}}$ ($\%$) metrics on the NUS datatset. The best performance is highlighted in bold.
Method Nature Human Activities Animals Technology Avg
Mcom $\uparrow$ Macc $\uparrow$ Mcom $\uparrow$ Macc $\uparrow$ Mcom $\uparrow$ Macc $\uparrow$ Mcom $\uparrow$ Macc $\uparrow$ Mcom $\uparrow$ Macc $\uparrow$
VQA BLIP-2 (Li et al., 2023) 70.76 60.54 70.52 59.94 71.88 59.99 71.48 60.74 71.16 60.30
InstructBLIP (Dai et al., 2023) 64.40 62.70 64.96 62.07 65.20 62.13 66.16 62.52 65.18 62.36
LLaVA (Liu et al., 2024c) 71.68 60.33 71.00 61.27 73.12 59.80 71.12 60.62 71.73 60.51
MiniGPT-4 (Zhu et al., 2024b) 76.76 49.97 77.32 50.30 70.36 49.15 69.84 50.30 73.57 49.93
Caption BLIP-2 (Li et al., 2023) 72.68 63.01 70.92 61.98 73.08 62.31 73.36 62.56 72.51 62.47
InstructBLIP (Dai et al., 2023) 72.24 65.73 72.88 66.15 72.76 64.98 72.12 65.57 72.50 65.61
LLaVA (Liu et al., 2024c) 74.76 61.71 73.32 63.77 75.08 61.21 72.44 63.80 73.90 62.62
MiniGPT-4 (Zhu et al., 2024b) 77.84 43.62 78.00 43.98 73.80 44.40 71.36 44.06 75.25 44.02
@@ 4.1. Experiment Setup
Table 3. Comparison of various setting for CoA in $\text{M}_{\text{com}}$ ($\%$) and $\text{M}_{\text{acc}}$ ($\%$) metrics on the VOC datatset. Due to page limitations, the Caption Action, Self-Correct Action, Appearance Action, Relationship Action and Final Action are abbreviated as Action 1 $\sim$ 5.
Type Vehicles Animals Indoor Outdoor Avg
Mcom $\uparrow$ Macc $\uparrow$ Mcom $\uparrow$ Macc $\uparrow$ Mcom $\uparrow$ Macc $\uparrow$ Mcom $\uparrow$ Macc $\uparrow$ Mcom $\uparrow$ Macc $\uparrow$
(i) LLaVA 67.65 74.64 67.38 76.89 75.15 75.54 65.77 74.27 68.99 75.34
@@
To evaluate the accuracy of each predicted label, we leverage the RAM model to score each predicted entity label  [^1] RAM is a large-scale multi-label model trained on extensive multi-label datasets, has demonstrated significant recognition accuracy in multi-label tasks (Zhang et al., 2024b).. Given a test image and a set of labels, RAM calculates a confidence score for each label in the list, filtering out labels that align with the image based on confidence thresholds. The accuracy score $AS_{i}$ of the predicted label list is then computed using the following formula:
$\small AS_{i}=\sum_{j}^{L}Sigmod(RAM(Y_{i}^{j})),$
where $RAM(Y_{i}^{j})$ denotes the confidence score calculated by RAM model for the label $j$ from predicted labels set $Y_{i}$, $L$ denotes the number of $Y_{i}$. A straightforward scoring rule is applied: one point is added for each correct prediction, and one point is subtracted for each incorrect prediction. Finally, we calculate the $\text{M}_{\text{acc}}=\sum_{i}^{N}AS_{i}/N$ to assess the accuracy of the predicted labels.
4.2. Quantitative and Qualitative Results
Comprehensive performance results for CoA and the baseline methods are detailed in Table 1 and Table 2 for both $\text{M}_{\text{acc}}$ ($\%$) and $\text{M}_{\text{com}}$ ($\%$). Our proposed method consistently surpasses all baselines across all evaluation metrics and datasets. Notably, (1) CoA exceeds the second-best method by an average of $11.56\%$ in $\text{M}_{\text{acc}}$ and $6.01\%$ in $\text{M}_{\text{com}}$ on the VOC dataset; (2) it demonstrates a $10.77\%$ average improvement in $\text{M}_{\text{acc}}$ and $6.21\%$ in $\text{M}_{\text{com}}$ on the COCO dataset; and (3) it surpasses the second-best method by $13.4\%$ in $\text{M}_{\text{acc}}$ and $4.45\%$ in $\text{M}_{\text{com}}$ on the NUS dataset. These results highlight the effectiveness of CoA, achieving the best performance both in terms of semantic comprehensiveness and accuracy of predicted labels.